# Márton Benedek 1122    Péter Biró 2211    Gergely Csáji 22 Matthew Johnson 33

### v2

@@ 4.6 Simulation Results for ${\mathbf{\ell=\infty}}$ and Comparison with ${\mathbf{\ell=2}}$
Influence of Using Credits. We now discuss the power of our credit system. Does it effectively reduce deviations from the target allocation? Can it keep deviations on a consistently low level in the long-term? Theoretically it is possible that credits keep accumulating, which would make them ineffective (see [10] for a theoretical example of this behaviour when $\ell=2$).
@@
In Figures 10 and 11 we show to what extent credits accumulate for equal and varying country sizes, respectively. If credits are not incorporated, we can still compute and track them as we did in these two figures. We note that over a period of 24 rounds, credits accumulate more and more (but at different rates) under all three scenarios arbitary, d1 and lexmin, even though especially under d1 and lexmin we do find solutions that are relatively close to the target allocations (as we saw from Figures 4 and 5). However, only when we also incorporate credits we not only find solutions that are close to the target solutions but that also ensure stability. We find that the Banzhaf* value under lexmin+c is also the best in maintaining consistently low levels of credits, much like in the case of $\ell=2$ [10]. Again, the main difference between $\ell=\infty$ and $\ell=2$ is that the arbitrary scenario performs much worse for $\ell=2$.
@@
Figure 10: Average credits for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
@@
Figure 11: Average credits for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
@@
Deviations for Worst-off Countries. Finally, we consider the ratio of the average maximum relative deviation and the average total relative deviation as a measure for the proportion of the total country deviations being due to the worst-off country. We call this ratio the relative ratio. For stability reasons, low relative ratios are preferable. From Table 3 we conclude that for the most interesting scenario lexmin+c and equal country sizes, the relative ratio decreases as the number of countries increases (as expected). In this figure we also included the relative ratios for $\ell=2$ (generated from the data used in [10]). We note that the relative ratio is slightly better for $\ell=2$. This extends to varying countries as well; see Appendix 0.C. In this appendix we also included the figures for the other scenarios as well. From these figures, we note that for d1+c the relative ratio for $\ell=\infty$ is lower than for $\ell=2$. The arbitrary scenario showed consistently lower deviations for $\ell=\infty$ than for $\ell=2$. However, this is not the case for the relative ratio. For equal country sizes, $\ell=\infty$ leads to consistently higher relative ratios than $\ell=2$ does, whereas for varying country set sizes the opposite holds. For the remaining scenarios (lexmin and d1), there is no clear distinction.
@@ 4.7 Computation Times
Table 4: Average CPU time for a single 24-round simulation instance, broken down into the different computational tasks. Here, the times for data preparation and graph building are average times taken over all scenarios and solution concepts, whereas the time for each solution concept is the average time taken over all scenarios. The total times for the scenarios are average times taken over all solutions concepts, where “total” refers to the total computation time, which includes computing the initial allocations, and only excludes the time for data preparation and graph building.
@@ 5 Conclusions
We introduced the class of partitioned permutation games and proved a number of complexity results that contrast known results for partitioned matching games. As open theoretical problems we ask: what is the complexity of computing the nucleolus for permutation games and for partitioned permutation games?
@@
Our new results guided our simulations for IKEPs with up to ten countries, with exchange bound $\ell=\infty$. Our simulations showed a significant improvement in the total number of kidney transplants over the case where $\ell=2$ [10] at the expense of long cycles. In our simulations, we first let all countries be of the same size. We then examined in the same way as done in [10], whether our conclusions would change for varying country sizes. As in [10], this turned out not to be the case.
@@
For future research we will consider the more realistic exchange bounds $\ell\in\{3,4,5\}$. We note that the simulations done in [12, 32] were only for $\ell=3$; a more limited number of solution concepts; for IKEPs with up to four countries; and only for scenarios that use weakly close optimal solutions. They also used different data sets. For our follow-up study for $\ell\in\{3,4,5\}$, we must now also overcome, as in [12, 32], the additional computational obstacle of not being able to compute an optimal solution for a compatibility graph in a KEP round and the values $v(S)$ of the associated permutation game in polynomial time (see Theorem 1.1). Current techniques for $\ell\in\{3,4,5\}$ therefore involve, besides ILPs based on the edge-formulation, ILPs based on the cycle-formulation, with a variable for each cycle of length at most $\ell$ (see, for example, [21, 23, 26]). Hence, computing a single value $v(S)$ will become significantly more expensive, and even more so for increasing $\ell$. For expensive solution concepts, such as the Shapley or nucleolus, we must compute an exponential number of values $v(S)$. Without any new methods, we expect it will not be possible to do this for simulations up to the same number of countries (ten) as we did for $\ell=\infty$.
Finally, we note that matching games and permutation games are usually defined on edge-weighted graphs. It is readily seen that all our positive theoretical results can be generalized to this setting. In kidney exchange, the primary goal is still to help as many patients as possible, but edge weights might be used to represent transplant utilities. Hence, it would also be interesting to do simulations in the presence of edge weights. We leave this for future research as well.
@@
Acknowledgments. Benedek was supported by the National Research, Development and Innovation Office of Hungary (OTKA Grant No. K138945); Biró by the Hungarian Scientific Research Fund (OTKA, Grant No. K143858) and the Hungarian Academy of Sciences (Momentum Grant No. LP2021-2); Csáji by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858, the ÚNKP-23-3 New National Excellence Program of the Ministry for Culture and Innovation and by the Doctoral Student Scholarship Program (number C2258525) of the Cooperative Doctoral Program of the Ministry of Innovation and Technology financed by the National Research, Development, and Innovation Fund; and Paulusma was supported by the Leverhulme Trust (Grant RF-2022-607) and EPSRC (Grant EP/X01357X/1).
This work has made use of the Hamilton HPC Service of Durham University. In particular, we thank Rob Powell for all his help with the initial set up of our simulations.
@@ Appendix 0.A Average Maximum Relative Deviations
Figure 12: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
Figure 13: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.

### v3

@@ 4.6 Simulation Results for $\mathbf{\ell=\infty}$} and Comparison with $\mathbf{\ell=2}$}
Influence of Using Credits. We now discuss the power of our credit system. Does it effectively reduce deviations from the target allocation? Can it keep deviations on a consistently low level in the long-term? Theoretically it is possible that credits keep accumulating, which would make them ineffective (see [BBPY24] for a theoretical example of this behaviour when $\ell=2$).
@@
In Figures 10 and 11 we show to what extent credits accumulate for equal and varying country sizes, respectively. If credits are not incorporated, we can still compute and track them as we did in these two figures. We note that over a period of 24 rounds, credits accumulate more and more (but at different rates) under all three scenarios arbitary, d1 and lexmin, even though especially under d1 and lexmin we do find solutions that are relatively close to the target allocations (as we saw from Figures 5 and 6). However, only when we also incorporate credits we not only find solutions that are close to the target solutions but that also ensure stability. We find that the Banzhaf* value under lexmin+c is also the best in maintaining consistently low levels of credits, much like in the case of $\ell=2$ [BBPY24]. Again, the main difference between $\ell=\infty$ and $\ell=2$ is that the arbitrary scenario performs much worse for $\ell=2$.
@@
Figure 10: Average credits for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
@@
Figure 11: Average credits for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries is $n=10$ and the period ranges from $1$ to $24$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which also behaved best for $\ell=2$.
@@
Deviations for Worst-off Countries. Finally, we consider the ratio of the average maximum relative deviation and the average total relative deviation as a measure for the proportion of the total country deviations being due to the worst-off country. We call this ratio the relative ratio. For stability reasons, low relative ratios are preferable. From Table 6 we conclude that for the most interesting scenario lexmin+c and equal country sizes, the relative ratio decreases as the number of countries increases (as expected). In this figure we also included the relative ratios for $\ell=2$ (generated from the data used in [BBPY24]). We note that the relative ratio is slightly better for $\ell=2$. This extends to varying countries as well; see Appendix 0.C. In this appendix we also included the figures for the other scenarios as well. From these figures, we note that for d1+c the relative ratio for $\ell=\infty$ is lower than for $\ell=2$. The arbitrary scenario showed consistently lower deviations for $\ell=\infty$ than for $\ell=2$. However, this is not the case for the relative ratio. For equal country sizes, $\ell=\infty$ leads to consistently higher relative ratios than $\ell=2$ does, whereas for varying country set sizes the opposite holds. For the remaining scenarios (lexmin and d1), there is no clear distinction.
@@ 4.7 Computation Times
Table 7: Average CPU time for a single 24-round simulation instance, broken down into the different computational tasks. Here, the times for data preparation and graph building are average times taken over all scenarios and solution concepts, whereas the time for each solution concept is the average time taken over all scenarios. The total times for the scenarios are average times taken over all solutions concepts, where “total” refers to the total computation time (using up to $32$ threads for each simulation instance), which includes computing the initial allocations, and only excludes the time for data preparation and graph building.
@@ 5 Conclusions
We introduced the class of partitioned permutation games and first proved a number of complexity results.
In particular, we showed that every partitioned permutation game has a nonempty core and that we can compute a core allocation of a partitioned permutation game in polynomial time. We also proved that for partitioned permutation games of fixed width $c$, the problem of deciding if an allocation is in the core is polynomial-time solvable if $c=1$ but is already coNP-complete if $c=2$. Moreover, we proved that in this setting even the problem of finding a weakly optimal solution for a given target allocation $x$ is NP-hard, and we gave an exact randomized XP-time algorithm for solving this problem. Our theoretical results contrast some known results for partitioned $\ell$-permutation games where $\ell$ is some constant, as surveyed in Table 1. As an open theoretical problem we ask:
Open Problem 1
Determine the complexity of computing the nucleolus for permutation games and for partitioned permutation games.
@@
Our new results guided our simulations for IKEPs, where we could include up to ten countries, with exchange bound $\ell=\infty$. Our simulations showed the clear benefits of using a credit system with strongly close optimal solutions, namely an improvement in balance up to 56%. The exact improvement depends on the choice of solution concept, with the Banzhaf* value yielding the best results, namely on average, a deviation of at most 0.90%. Our simulations showed a significant improvement (46% on average) in the total number of kidney transplants if cycles of any length are allowed compared to the case where $\ell=2$ [BBPY24]. In our simulations, we first let all countries be of the same size. We then examined, in the same way as was done in [BBPY24], whether our conclusions would change for varying country sizes. Just as in [BBPY24], this turned out not to be the case.
For future research, it would be interesting to determine to what extent our theoretical and experimental results change if we allow non-directed (altruistic) donors or multiple donors registering for one patient. We did not consider such types of donors for this paper, as our goal was to compare our results with the previously known results for $\ell=2$, where non-directed donors and multiple donors for one patient were also excluded. Moreover, we recall that matching games and permutation games may be defined on edge-weighted graphs, in which each edge $e\in A$ has some weight $w(e)$ reflecting the expected utility of the corresponding kidney transplant. It would be interesting to do simulations in the presence of edge weights. Furthermore, we recall that more complex hierarchical optimization criteria than only optimizing the number of transplants are used in European kidney exchange [Bi_etal2019, Bi_etal2021]. We leave conducting simulations with such criteria for future research as well.
@@
Finally, for future research we will consider the more realistic exchange bounds $\ell\in\{3,4,5\}$. We note that the simulations done in [BGKPPV20, KNPV20] were only for $\ell=3$; a more limited number of solution concepts; for IKEPs with up to four countries; and only for scenarios that use weakly close optimal solutions. They also used different data sets. For our follow-up study for $\ell\in\{3,4,5\}$, we must now also overcome, as in [BGKPPV20, KNPV20], the additional computational obstacle of not being able to compute an optimal solution for a compatibility graph in a KEP round and the values $v(S)$ of the associated permutation game in polynomial time (see Theorem 1.1). Current techniques for $\ell\in\{3,4,5\}$ therefore involve, besides ILPs based on the edge-formulation, ILPs based on the cycle-formulation, with a variable for each cycle of length at most $\ell$ (see, for example, [CKVR13, DGGKMP, DMPST16]). Hence, computing a single value $v(S)$ will become significantly more expensive, and even more so for increasing $\ell$. For expensive solution concepts, such as the Shapley or nucleolus, we must compute an exponential number of values $v(S)$. Without any new methods, we expect it will not be possible to do this for simulations up to the same number of countries (ten) as we did for $\ell=\infty$.
@@
Acknowledgments. Benedek was supported by the National Research, Development and Innovation Office of Hungary (OTKA Grant No. K138945); Biró by the Hungarian Scientific Research Fund (OTKA, Grant No. K143858) and the Hungarian Academy of Sciences (Momentum Grant No. LP2021-2); Csáji by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858, and by the Doctoral Student Scholarship Program (number C2258525) of the Cooperative Doctoral Program of the Ministry of Innovation and Technology financed by the National Research, Development, and Innovation Fund; and Paulusma was supported by the Leverhulme Trust (Grant RF-2022-607) and EPSRC (Grant EP/X01357X/1). This work has made use of the Hamilton HPC Service of Durham University. In particular, we thank Rob Powell for all his help with the initial set up of our simulations.
We also thank the three anonymous reviewers of our paper for all their useful comments for improving the readability of our paper.
@@ Appendix 0.A Average Maximum Relative Deviations
Figure 12: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
Figure 13: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [BBPY24] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
@@ Appendix 0.B Additional Information on Incomplete Simulations
Tables 8 and 9 give complete breakdowns of the number of incomplete simullations instances for equal country sizes and varying country sizes, respectively. Figures 16 and 17 are analogous to Figure 5 and 6, respectively; the only difference is that for Figures 16 and 17 we left out the incomplete instances.
@@ Appendix 0.C Relative Ratios
Recall that the relative ratio is the ratio of the average maximum relative deviation and the average total relative deviation. In Table 6 we showed the relative ratios for equal country sizes under lexmin+c both for $\ell=\infty$ and $\ell=2$, and we also showed their difference. In Tables 1018 we show the same information for the other scenarios and also for varying country sizes (where the relative ratios for $\ell=2$ where generated from the data used in [BBPY24]).
@@
Appendix 0.D Full Results on Number of Kidney Transplants
Tables 4 and 5 only showed results for the average number of kidney transplants for $n=10$. We included the information from these two tables for $\ell=\infty$ in Tables 19 and 20, together with the results for $n=4,\ldots,9$, for the case where $\ell=\infty$; we refer to [BBPY24] for the results for $n=4,\ldots,9$ for the case where $\ell=2$.
Solution concepts/n 4 5 6 7 8 9 10
benefit value arbitrary 1781.89 1780.60 1775.74 1771.30 1767.59 1763.44 1781.58
d1 1782.15 1781.01 1775.31 1771.21 1767.38 1763.04 1781.65
d1+c 1781.52 1780.01 1774.51 1770.22 1766.34 1762.37 1781.46
lexmin 1782.25 1781.25 1775.47 1771.56 1767.87 1763.19 1782.58
lexmin+c 1781.69 1780.12 1774.99 1770.71 1766.30 1761.76 1780.31
contribution value arbitrary 1781.89 1780.60 1775.74 1771.30 1767.59 1763.44 1781.58
d1 1782.70 1781.30 1775.88 1772.07 1768.31 1763.31 1781.52
d1+c 1781.43 1780.41 1775.43 1770.90 1766.73 1762.68 1780.50
lexmin 1782.20 1781.42 1775.86 1771.93 1768.04 1763.31 1782.52
lexmin+c 1781.80 1780.58 1774.87 1770.21 1766.24 1762.30 1780.05
Nucleolus arbitrary 1781.89 1780.60 1775.67 1771.43 1767.64 1763.17 1781.58
d1 1781.79 1780.99 1775.76 1771.53 1767.26 1762.97 1781.80
d1+c 1781.24 1779.79 1775.30 1770.68 1767.00 1762.30 1780.45
lexmin 1782.02 1781.11 1775.30 1771.99 1767.42 1763.19 1781.98
lexmin+c 1781.43 1779.96 1774.86 1770.74 1765.81 1761.97 1780.12
Shapley value arbitrary 1781.89 1780.60 1775.78 1771.38 1767.57 1763.14 1781.58
d1 1782.25 1781.22 1775.81 1771.75 1767.06 1763.34 1781.54
d1+c 1781.67 1780.88 1775.45 1771.11 1766.68 1762.22 1780.92
lexmin 1782.43 1781.90 1775.38 1771.73 1767.90 1763.37 1782.02
lexmin+c 1781.61 1780.79 1775.50 1771.13 1767.13 1762.65 1780.57
Banzhaf value arbitrary 1781.89 1780.60 1775.71 1771.46 1767.50 1763.31 1781.58
d1 1782.14 1781.14 1775.34 1771.91 1767.57 1763.15 1781.83
d1+c 1781.67 1780.33 1774.87 1771.37 1766.51 1762.34 1780.38
lexmin 1782.15 1781.15 1775.62 1771.97 1767.99 1763.16 1782.46
lexmin+c 1781.96 1780.58 1774.87 1771.08 1766.94 1762.79 1780.66
Banzhaf* value arbitrary 1781.89 1780.60 1775.71 1771.46 1767.50 1763.31 1781.58
d1 1782.14 1781.14 1775.34 1771.91 1767.57 1763.15 1781.83
d1+c 1781.77 1780.03 1774.83 1771.38 1767.00 1762.71 1781.08
lexmin 1782.15 1781.15 1775.62 1771.97 1767.99 1763.16 1782.46
lexmin+c 1782.04 1781.03 1774.78 1771.32 1767.24 1762.70 1781.24
Tau value arbitrary 1781.89 1780.60 1775.65 1771.36 1767.34 1763.62 1781.58
d1 1781.87 1780.67 1775.14 1771.22 1767.35 1762.70 1782.22
d1+c 1781.80 1780.08 1774.28 1770.34 1766.54 1761.85 1780.90
lexmin 1782.41 1780.63 1775.40 1771.60 1767.60 1763.42 1782.12
lexmin+c 1781.13 1780.01 1774.42 1770.97 1766.31 1761.81 1780.45
Table 19: Average number of kidney transplants for equal country sizes for $n\in\{4,\ldots,10\}$ when $\ell=\infty$.
Solution concepts/n 4 5 6 7 8 9 10
benefit value arbitrary 1775.56 1780.86 1765.16 1745.57 1754.08 1731.64 1763.57
d1 1775.07 1781.09 1765.15 1745.78 1753.74 1731.19 1763.39
d1+c 1774.24 1780.35 1763.89 1745.75 1753.46 1731.12 1763.37
lexmin 1775.51 1781.39 1764.70 1745.66 1753.83 1731.51 1763.73
lexmin+c 1774.38 1780.40 1763.87 1744.95 1752.83 1730.72 1762.45
contribution value arbitrary 1775.56 1780.86 1765.16 1745.57 1754.08 1731.64 1763.57
d1 1775.17 1780.39 1765.10 1746.20 1753.91 1731.31 1763.45
d1+c 1774.09 1780.19 1763.82 1745.31 1753.04 1730.92 1762.90
lexmin 1775.52 1780.88 1765.52 1746.61 1754.36 1732.38 1763.80
lexmin+c 1773.54 1778.83 1763.59 1744.50 1752.82 1730.44 1761.84
Nucleolus arbitrary 1774.12 1781.25 1764.39 1746.27 1753.55 1731.72 1763.66
d1 1774.18 1780.85 1764.08 1746.64 1753.38 1731.13 1764.02
d1+c 1773.08 1779.98 1763.28 1745.61 1752.75 1730.71 1762.79
lexmin 1774.23 1781.34 1764.32 1746.04 1753.41 1731.24 1763.74
lexmin+c 1772.92 1780.06 1762.92 1745.32 1751.82 1730.22 1762.19
Shapley value arbitrary 1774.02 1781.25 1764.42 1746.17 1753.44 1731.64 1764.04
d1 1774.38 1781.17 1763.92 1746.27 1753.60 1731.43 1763.09
d1+c 1773.47 1779.94 1763.70 1745.30 1753.14 1731.43 1762.99
lexmin 1774.24 1781.50 1764.67 1745.89 1753.84 1731.73 1763.71
lexmin+c 1773.51 1780.54 1764.09 1745.53 1752.84 1730.54 1762.60
Banzhaf value arbitrary 1775.90 1780.86 1765.00 1745.64 1754.25 1731.39 1763.24
d1 1775.17 1781.00 1764.38 1746.30 1754.43 1731.38 1763.34
d1+c 1775.25 1780.41 1764.60 1745.27 1753.50 1731.05 1762.85
lexmin 1775.43 1781.02 1765.33 1745.74 1754.33 1732.17 1763.47
lexmin+c 1774.90 1780.01 1764.23 1745.56 1753.40 1731.02 1762.85
Banzhaf* value arbitrary 1775.90 1780.86 1765.00 1745.64 1754.25 1731.39 1763.24
d1 1775.17 1781.00 1764.38 1746.30 1754.43 1731.38 1763.34
d1+c 1774.62 1780.69 1764.32 1745.91 1753.87 1731.27 1763.09
lexmin 1775.43 1781.02 1765.33 1745.74 1754.33 1732.17 1763.47
lexmin+c 1775.02 1780.43 1764.81 1745.63 1753.82 1731.49 1762.97
Tau value arbitrary 1775.46 1780.86 1765.12 1745.75 1754.50 1731.55 1763.27
d1 1774.72 1780.85 1764.86 1745.48 1753.78 1731.42 1763.54
d1+c 1774.50 1779.87 1763.82 1746.00 1753.06 1731.11 1762.99
lexmin 1774.90 1781.11 1764.47 1745.77 1754.02 1731.48 1763.62
lexmin+c 1774.46 1779.77 1763.52 1745.32 1753.29 1730.23 1761.84
Table 20: Average number of kidney transplants for varying country sizes for $n\in\{4,\ldots,10\}$ when $\ell=\infty$.